OpenStudy (anonymous):

What is the first step in solving 4(x + 2) = 10 + x?

OpenStudy (theviper):

solve brackets :)

OpenStudy (anonymous):

4x+8=10+x

OpenStudy (theviper):

\[\LARGE{4(x + 2) = 10 + x}\]\[\Large{4x+8=10+x}\]\[\Large{4x-x=10-8}\]\[\Large{3x=2}\]\[\Large{x=\frac{2}{3}}\]

OpenStudy (theviper):

\[\Huge{\color{red}{\checkmark}}\]